# The equational theory of the Weihrauch lattice with multiplication

### v1

@@ Abstract
We study the equational theory of the Weihrauch lattice with multiplication, meaning the collection of equations between terms built from variables, the lattice operations $\sqcup$,$\sqcap$ , ⊓, the product $\times$, and the finite parallelization $*{}^{*}$ which are true however we substitute Weihrauch degrees for the variables. We provide a combinatorial description of these in terms of a reducibility between finite graphs, and moreover, show that deciding which equations are true in this sense is complete for the third level of the polynomial hierarchy.
@@ 1 Introduction
The Weihrauch degrees $\mathfrak{W}$ come with a rich algebraic structure. Here, we consider the lattice operations $\sqcup$,$\sqcap$ , ⊓, the product $\times$ and the finite parallelization $*{}^{*}$. They were the first operations on Weihrauch degrees studied in the literature [3, 2, 14]. In order to better understand the structure of the Weihrauch degrees, we would like to characterize its equational theory, i.e. we want to identify which equations between terms over the signature $(\mathfrak{W},\sqcap,\sqcup,\times,1,(-)^{*})$ are true for every instantiation of the variables by Weihrauch degrees. It was already observed in [5] that the equational theory of $(\mathfrak{W},\sqcap,\sqcup)$ is the theory of distributive lattices, as $(\mathfrak{W},\sqcap,\sqcup)$ is a distributive lattice itself, and every countable distributive lattice embeds into $(\mathfrak{W},\sqcap,\sqcup)$ (via the Medvedev degrees).
@@
We start by investigating the equational theory of $(\mathfrak{W},\sqcap,\times)$. While we did not succeed in identifying a finite axiomatization for it (and in fact, conjecture that there might be none using only equations), we provide a combinatorial characterization in terms of reductions between finite graphs. This in turn will allow us to show that determining universal validity of equations in $(\mathfrak{W},\sqcap,\times)$ is $\Sigma^{p}_{2}$-complete. Since we only work with substructures that include $\sqcap$, we can write inequalities instead of equations by treating $t\leq u$ as an abbreviation for $t=t\sqcap u$.
@@
We are not aware of axioms systems matching exactly those we provided for $(\mathfrak{W}_{\bullet},\sqcap,\times)$ in the literature. Despite of this, there are tentative connections to be made between notions of correctness in substructural logics being phrased in terms of graphs and our notion of reduction in terms of graphs. In particular, [6] has a strikingly similar notion of reducibility between graphs generalizing boolean formulas.
Although our notion of reducibility is more complex, this seems to indicate casting our axioms or our general notion of graph reducibility in a deep inference system might be of relevance in uncovering a complete axiomatization for the theories we consider.
@@ 2 Introducing the operations
Additional axiom for the pointed case $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$ $\begin{array}[]{cr}1\leq_{\bullet}a&\text{bottom element}\\ \end{array}$
@@
The proposed axiomatization in Figure 1 does not capture all valid inequalities in $\mathfrak{W}$.
Proposition 4.
Figure 1 is not a complete axiomatization of inequality over $(\mathfrak{W},\sqcap,\times)$.
Proof.
An inequality which holds in $(\mathfrak{W},\sqcap,\times)$ but cannot be derived from the axioms in Figure 1 is:
$a\times(b\sqcap c)\leq a\times(b\sqcap(a\times c))$
The reduction works by copying the input to $a$ on the left to both occurrences of $a$ on right, and sending the inputs for $b$ and $c$ on the left to their right-side counterparts. The outer reduction witness ignores the response from the second $a$ if applicable, and copies every other answer to $a$ and $b$ or $c$ on the left.
To see that the reduction cannot be derived from the axioms, we start with the left-hand side and try to rewrite the term. Applying monotonicity or relevance yields terms that are no longer reducible to the term on the right. Applying weakening to $b\sqcap c$ also destroys the reduction. Half-distributivity does not apply. Thus, there is no potential first step for rewriting the left-hand expression to the right-hand expression. ∎
Note that $a\times(b\sqcap c)\leq_{\bullet}a\times(b\sqcap(a\times c))$ is derivable from the axioms in Figure 1, as here we can use $c\leq_{\bullet}(a\times c)$ together with monotonicity to immediately derive the right from the left.
@@
There is no way of extending the axiomatization of inequality over $(\mathfrak{W},\sqcap,\times,1)$ with finitely many inequalities to obtain a complete system.
We were unable to find a true inequality in $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$ that does not follow from the stated axioms, and thus offer up the following:
Conjecture 6.
$(\mathfrak{W}_{\bullet},\sqcap,\times,1)$ is fully axiomatized by Figure 1.
@@ 3 A combinatorial notion of reducibility for terms
The graphs that arise as $\llbracket u\rrbracket$ for some term $u$ are known as cographs. They can also be characterized as those graphs that do not have an induced subgraph forming a line of length $3$ [7].
@@
Extending our results for inequalities $t\leq_{\bullet}u$ built from ,$\sqcap$,$\times$ , × to terms also include the constant $1$ is straightforward, as our axioms for the pointed degrees allow us to simplify any term over (,×$(\sqcap,\times,1)$ to produce either just $1$ or a term over (,×)square-intersection($\sqcap$,$\times$)( ⊓ , × ).
For inequalities $t\leq u$ it is less straightforward. We can extend our interpretations of terms as coloured cographs to include $1$ by mapping $1$ to an uncoloured vertex. In the definition of a reduction map the uncoloured vertices are ignored other than in defining what the maximal cliques are. This means that the reduction map does not have to be defined on the uncoloured vertices, and the image of a maximal clique in $\llbracket u\rrbracket$ has to contain – up to some uncoloured vertices – a maximal clique in $\llbracket t\rrbracket$.
@@ 4 Introducing joins and parallelization
Parallelization $\begin{array}[]{cr}a\leq a^{*}\qquad(a^{*})^{*}\leq a^{*}&\text{closure operator}\\ a^{*}\times a^{*}\leq a^{*}&\text{duplicable}\\ (a\sqcup b)^{*}=a^{*}\times b^{*}&\text{interaction with $\sqcup$}\\ (a\sqcap b)^{*}=a^{*}\sqcap b^{*}&\text{interaction with $\sqcap$}\\ (a\times b)^{*}=1\sqcup a\times a^{*}\times b\times b^{*}&\text{interaction with $\times$}\\ 1^{*}=1&\text{interaction with $1$}\\ \end{array}$
@@
However, we can re-write any term in linear time into one where parallelization is applied only to variables. This is obvious in the pointed case, since the rule $(a\times b)^{*}=1\sqcup a\times a^{*}\times b\times b^{*}$ together with pointedness yields $(a\times b)^{*}=a^{*}\times b^{*}$. In general, an easy induction shows that any term of the form $*t=\left(\bigtimes_{i=1}^{n}t_{i}\right)^{*}$ is derivably equivalent to $1\sqcup\bigtimes_{i=1}^{n}\left(t_{i}\times t_{i}^{*}\right)$ via a linear-size derivation.
@@ 5 Complexity of deciding validity
The mapping $f$ exists essentially because $K$ is chosen large enough to ensure that we can cover all occurrences of variables in $F$ as described above. It is quite straightforward to check that $f$ is a combinatorial reduction.
@@ Acknowledgments
We are grateful to Manlio Valenti and Lê Thành Dũng Nguy~n for discussions, and also to Matteo Acclavio for some pointers to the literature.

### v2

@@ Abstract
We study the equational theory of the Weihrauch lattice with multiplication, meaning the collection of equations between terms built from variables, the lattice operations $\sqcup$,$\sqcap$ , ⊓, the product $\times$, and the finite parallelization which are true however we substitute Weihrauch degrees for the variables. We provide a combinatorial description of these in terms of a reducibility between finite graphs, and moreover, show that deciding which equations are true in this sense is complete for the third level of the polynomial hierarchy.
Erratum
The first version of this preprint claimed that our theory was not complete because $a\times(b\sqcap c)\leq a\times(b\sqcap(c\times a))$ was not derivable without the pointedness axiom. This is false. We have removed the erroneous statement and added a proof that the axiomatization we propose is complete for Weihrauch degrees if and only if it is complete for pointed degrees if $1\leq a$ is added (Theorem 15).
@@ 1 Introduction
The Weihrauch degrees $\mathfrak{W}$ come with a rich algebraic structure. Here, we consider the lattice operations $\sqcup$,$\sqcap$ , ⊓, the product $\times$ and the finite parallelization . They were the first operations on Weihrauch degrees studied in the literature [3, 2, 14]. In order to better understand the structure of the Weihrauch degrees, we would like to characterize its equational theory, i.e. we want to identify which equations between terms over the signature $(\mathfrak{W},\sqcap,\sqcup,\times,1,(-)^{*})$ are true for every instantiation of the variables by Weihrauch degrees. It was already observed in [5] that the equational theory of $(\mathfrak{W},\sqcap,\sqcup)$ is the theory of distributive lattices, as $(\mathfrak{W},\sqcap,\sqcup)$ is a distributive lattice itself, and every countable distributive lattice embeds into $(\mathfrak{W},\sqcap,\sqcup)$ (via the Medvedev degrees).
@@
We start by investigating the equational theory of $(\mathfrak{W},\sqcap,\times)$. While we only conjecture that the axiomatization we propose for it is complete, we provide a combinatorial characterization in terms of reductions between finite graphs. This in turn will allow us to show that determining universal validity of equations in $(\mathfrak{W},\sqcap,\times)$ is $\Sigma^{p}_{2}$-complete. Since we only work with substructures that include $\sqcap$, we can write inequalities instead of equations by treating $t\leq u$ as an abbreviation for $t=t\sqcap u$.
@@
We are not aware of axioms systems matching exactly those we provided for $(\mathfrak{W}_{\bullet},\sqcap,\times)$ in the literature. Despite of this, there are tentative connections to be made between notions of correctness in substructural logics being phrased in terms of graphs and our notion of reduction in terms of graphs. In particular, [6] has a strikingly similar notion of reducibility between graphs generalizing boolean formulas. this seems to indicate casting our axioms or our general notion of graph reducibility in a deep inference system might be of relevance in uncovering a complete axiomatization for the theories we consider.
@@ 2 Introducing the operations
Additional axiom for the pointed case $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$mdframed}
$\begin{array}[]{cr}1\leq_{\bullet}a&\text{bottom element}\\ \end{array}$
@@
We were unable to find a true inequality in $(\mathfrak{W},\sqcap,\times,1)$ that does not follow from the stated axioms, and thus offer up the following:
@@
$(\mathfrak{W},\sqcap,\times,1)$ is fully axiomatized by Figure 1.
We also show later that the axiomatization is complete for $\mathfrak{W}$ if and only if it the pointed version is complete for $\mathfrak{W}_{\bullet}$ (Theorem 15).
@@ 3 A combinatorial notion of reducibility for terms
The graphs that arise as $\llbracket u\rrbracket$ for some term $u$ are known as cographs. They can also be characterized as those graphs that do not have an induced subgraph forming a line of length $3$ [7]. Below we give some examples of terms and corresponding cographs.
The intuition is that a coloured graph $(V,E,c)$ can be read as the following Weihrauch problem, assuming the colours correspond themselves to Weihrauch problems: given for each vertex $v\in V$ an input $i_{v}$ to $c(v)$, output a maximal clique $C$ of $(V,E)$ together with solutions for every $i_{v}$s with $v\in C$. Under this interpretation, the problems denoted $\llbracket t\rrbracket$ and $t$ are Weihrauch-equivalent. This leads naturally to the following notion of combinatorial reduction between coloured graphs.
@@
Let us illustrate the notion of combinatorial reduction on our examples.
In the two examples above, we have a strict reduction. We do not specify the total maps between vertices as they are uniquely determined by the colors in these particular cases. The shaded areas correspond to maximal cliques in the graphs and they have matching colors when the reductions relate them. There are no reductions the other ways around: for the first scenario, this is because the clique $\{a,c\}$ on the left can only be mapped to a strict subclique of the triangular clique on the right by a color-preserving map. For the second case, where we have a unique color preserving-map, it is because $\{a,b\}$ does not contain a maximal clique in the rightmost graph.
@@
Extending our results for inequalities $t\leq_{\bullet}u$ built from ,$\sqcap$,$\times$ , × to terms also including the constant $1$ is straightforward, as our axioms for the pointed degrees allow us to simplify any term over (,×$(\sqcap,\times,1)$ to produce either just $1$ or a term over (,×)square-intersection($\sqcap$,$\times$)( ⊓ , × ).
For inequalities $t\leq u$, this is less straightforward. We can extend our interpretations of terms as coloured cographs to include $1$ by mapping $1$ to an uncoloured vertex. In the definition of a reduction map the uncoloured vertices are ignored other than in defining what the maximal cliques are. This means that the reduction map does not have to be defined on the uncoloured vertices, and the image of a maximal clique in $\llbracket u\rrbracket$ has to contain – up to some uncoloured vertices – a maximal clique in $\llbracket t\rrbracket$.
@@
4 Completeness does not depend on pointedness
Theorem 15.
The axioms of Figure 1 are complete for $(\mathfrak{W},\sqcap,\times,1)$ if and only if they are complete for $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$ when adding $1\leq_{\bullet}a$
To prove this theorem, we first establish some results for derivability that allows us to relate the pointed and non-pointed theories. For any finite set $X$ of variables, call $i_{X}$ the formula $1\sqcap\bigsqcap_{x\in X}x$. The intuition is that the corresponding Weihrauch problem requires at least an input for every $x$ to be provided, but does not answer any of the corresponding questions.
Lemma 16.
For any finite sets of variables $X$ and $Y$, we can derive $i_{X}\times i_{Y}=i_{X\cup Y}$.
Proof.
Firstly we can show that we have $i_{X}\times i_{X}=i_{X}$. This is because we have the relevance axiom and $i_{X}\times i_{X}\leq i_{X}\times 1=i_{X}$, since $i_{X}\leq 1$ by definition. Then, note that $i_{X^{\prime}}\leq i_{X}$ whenever $X\subseteq X^{\prime}$. So we have $i_{X\cup Y}=i_{X\cup Y}\times i_{X\cup Y}\leq i_{X}\times i_{Y}$. For the converse direction, it suffices to show that $i_{X}\times i_{Y}\leq 1$, $i_{X}\times i_{Y}\leq x$ for all $x\in X$ and, symmetrically, $i_{X}\times i_{Y}\leq y$ for all $y\in Y$. This is straightforward as $i_{X}\leq 1$ and $i_{X}\leq x$ for every $x\in X$. ∎
Lemma 17.
For any term $t$ whose set of free variables is $X$, we can derive $i_{X}\times t=t$.
Proof.
Deriving $t\leq i_{X}\times t$ is easy since $1\leq i_{X}$. For the converse, we proceed by induction over $t$. If $t=1$, this amounts to $1\times 1=1$. If $t=t_{1}\times t_{2}$, call $X_{1}$ and $X_{2}$ the free variables of $t_{1}$ and $t_{2}$ respectively. In this case we have $X=X_{1}\cup X_{2}$, so we can derive $i_{X}\times t\leq i_{X_{1}}\times t_{1}\times i_{X_{2}}\times t_{2}$ using Lemma 16 and then conclude using the inductive hypotheses and the monotonicity axiom for $\times$. If we have $t=t_{1}\sqcap t_{2}$, adopting similar naming conventions, we also have $i_{X}\times t\leq(i_{X_{1}}\times t_{1})\sqcap(i_{X_{2}}\times t_{2})$ using the fact that $i_{X}\leq i_{Y}$ for any $Y\subseteq X$, and we similarly using monotonicity of $\sqcap$. ∎
Lemma 18.
For any terms $t$ and $u$ such that all the free variables of $t$ and $u$ are contained in a finite set $X$, then we have that $i_{X}\times t\leq u$ is derivable if and only if $t\leq_{\bullet}u$ is derivable.
Proof.
The left-to-right direction is straightforward, as we have $1\leq_{\bullet}x$ for all $x\in X$ and therefore $1\leq_{\bullet}i_{X}$. For the right-to-left direction, we need to reason by induction on the size of a derivation $t\leq_{\bullet}u$ and make a case analysis on the last axiom applied.
  • If the last axiom applied is reflexivity and $t=u$, we can simply apply Lemma 17 since $X$ is a superset of the free variables of $t$.
  • If the last axiom applied is transitivity, i.e. we have two subderivations of $t\leq_{\bullet}v$ and $v\leq_{\bullet}u$ for some $v$, we can assume, without loss of generality, that all variables occurring in $v$ occur in $X$. Because if we had such a variable $a$ in $v$, we could get derivations of the same size of $t=t[1/a]\leq_{\bullet}v[1/a]$ and $v[1/a]\leq_{\bullet}u[1/a]=u$. Therefore, we can apply the induction hypothesis to get derivations of $i_{X}\times u\leq v$ and $i_{X}\times v\leq u$ which we can chain as follows
    $i_{X}\times t\leq i_{X}\times(i_{X}\times t)\leq i_{X}\times v\leq u$
  • If the last axiom applied is monotonicity of of $\times$, i.e. we have derivations $t_{1}\leq_{\bullet}u_{1}$, $t_{2}\leq_{\bullet}u_{2}$ and $t=t_{1}\times t_{2}$ and $u=u_{1}\times u_{2}$, then we can apply the inductive hypotheses to obtain $i_{X}\times t_{j}\leq u_{j}$ for $j\in\{1,2\}$ and conclude by
    $i_{X}\times t_{1}\times t_{2}\leq i_{X}\times t_{1}\times i_{X}\times t_{2}\leq u_{1}\times u_{2}$
  • Monotonicity of $\sqcap$ is handled similarly.
  • All of the other axioms shared between the pointed and non-pointed cases are handled trivially. So only $1\leq_{\bullet}t$ remains, for which it suffices to prove that $i_{X}\leq t$. This is done via an easy induction over $t$.
Proof of Theorem 15.
First let us assume that our theory is complete for $\mathfrak{W}_{\bullet}$ and that we have that $t\leq u$ combinatorially valid in $\mathfrak{W}$. In particular, our theory derives $t\leq_{\bullet}u$. Via combinatorial validity, we also know that means that the free variables of $u$ are also free variables of $t$; call $X$ the set of free variables of $t$. By Lemma 18, we have that our theory derives $i_{X}\times t\leq u$, but we also have by Lemma 17 that $t\leq i_{X}\times t$, so we can conclude that the theory is also complete for $\mathfrak{W}$.
Conversely, if the theory is complete for $\mathfrak{W}$ and $t\leq_{\bullet}u$ is combinatorially valid in $\mathfrak{W}_{\bullet}$, then, if $X$ is the set of free variables of $u$, $i_{X}\times t\leq u$ is easily seen to be combinatorially valid; the partial map given by the pointed reduction can be made total by mapping unmapped vertices to those in $\llbracket i_{X}\rrbracket$. Then, by completeness we have $i_{X}\times t\leq u$ derivable, and since $1\leq_{\bullet}i_{X}$, we a fortiori have $t=1\times t\leq_{\bullet}i_{X}\times t\leq u$ derivable. ∎
@@ 5 Introducing joins and parallelization
Parallelization {mdframed}
$\begin{array}[]{cr}a\leq a^{*}\qquad(a^{*})^{*}\leq a^{*}&\text{closure operator}\\ a^{*}\times a^{*}\leq a^{*}&\text{duplicable}\\ (a\sqcup b)^{*}=a^{*}\times b^{*}&\text{interaction with $\sqcup$}\\ (a\sqcap b)^{*}=a^{*}\sqcap b^{*}&\text{interaction with $\sqcap$}\\ (a\times b)^{*}=1\sqcup a\times a^{*}\times b\times b^{*}&\text{interaction with $\times$}\\ 1^{*}=1&\text{interaction with $1$}\\ \end{array}$
@@
However, we can re-write any term in linear time into one where parallelization is applied only to variables. This is obvious in the pointed case, since the rule $(a\times b)^{*}=1\sqcup a\times a^{*}\times b\times b^{*}$ together with pointedness yields $(a\times b)^{*}=a^{*}\times b^{*}$. In general, an easy induction shows that any term of the form $t=\left(\bigtimes_{i=1}^{n}t_{i}\right)^{*}$ is derivably equivalent to $1\sqcup\bigtimes_{i=1}^{n}\left(t_{i}\times t_{i}^{*}\right)$ via a linear-size derivation.
@@ 6 Complexity of deciding validity
The mapping $f$ exists essentially because $K$ is chosen large enough to ensure that we can cover all occurrences of variables in $F$ as described above.
Now, we claim that $f$ is a combinatorial reduction from $E_{L}\times F$ to $E_{R}\times A$. Let $C$ be maximal clique in $E_{R}\times A$. Then $C$ is of the form $C=E_{R}\times C^{\prime}$, where $C^{\prime}$ is a maximal clique in $A$. Further, $C^{\prime}$ is of the form
$C^{\prime}=\bigtimes\limits_{m=1}^{n}z_{m}^{K}$
where $z_{m}\in\{y_{m},\overline{y_{m}}\}$. Let $\theta\colon\{y_{1},\dots,y_{n}\}\to\{0,1\}$ be the valuation that sends $y_{j}$ to $1$ if $z_{j}=y_{j}$ and to 00 if $z_{j}=\overline{y_{j}}$. Then, by assumption, some conjunction of literals $\bigwedge_{i\in I}\ell_{j(i)}$ holds true under the valuations $\rho$ and $\theta$. Consider the corresponding maximal clique ×$\bigtimes_{i\in I}\ell_{j(i)}$ in $F$. Let $E_{L}^{\prime}$ be the maximal clique ×$\bigtimes_{m=1}^{k}w_{m}^{K}$ of $E_{L}$ where $w_{m}\in\{x_{m},\overline{x_{m}}\}$ and $w_{m}=x_{m}$ if and only if $\rho(x_{m})=0$. Then $E_{L}^{\prime}\times\bigtimes_{i\in I}\ell_{j(i)}$ is a maximal clique in $E_{L}\times F$. By construction, every vertex in $E_{L}$ is the image of a vertex in $E_{R}$ under $f$.
Now, by assumption, $\ell_{j(i)}$ is consistent with $\rho$ and $\theta$ in the sense that if $\ell_{j(i)}\in\{x_{m},\overline{x_{m}}\}$, then $\ell_{j}(i)=x_{m}$ if and only if $\rho(x_{m})=1$ and similarly for $\theta$ and $\ell_{j(i)}\in\{y_{m},\overline{y_{m}}\}$. But this implies by construction that every $\ell_{j(i)}\in\{y_{m},\overline{y}_{m}\}$ is the image of a vertex in $C^{\prime}$ under $f$ and every $\ell_{j(i)}\in\{x_{m},\overline{x}_{m}\}$ is the image of a vertex in $E_{R}$ under $f$. This shows that $f$ is a combinatorial reduction.
@@ Acknowledgments
We are grateful to Manlio Valenti and Lê Thành Dũng Nguyễn for discussions, and also to Matteo Acclavio for some pointers to the literature.